# SelectLLM: Can LLMs Select Important Instructions to Annotate?

### 20 Feb 2024 v2

@@ 1 Introduction
As it is relatively easy to access unlabelled instructions or texts from various sources, one promising solution to address these challenges is developing a way to select important unlabelled instructions for annotation, similar to active learning (Settles, 2009); from this, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. However, existing selection methods present limitations. Active learning-based algorithms focusing on density in the embedding space (Sener and Savarese, 2018) often overlook the complexity of instruction samples. Alternatively, output-based approaches that assess uncertainty in model predictions (Kung et al., 2023) grapple with high computational costs and the challenges posed by synthetic label quality (Figure 1). In light of these challenges, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we investigate the potential of LLMs to select effective unlabelled instructions, leveraging their vast knowledge base to discern the complexity and utility of each instruction without generating inferences.
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. At a high level, our method uses LLMs to estimate the usefulness and impact of each instruction without the corresponding labels; through the initial experiments presented in Figure 3, we first verify that LLMs do possess such a capability. To improve instruction selection via LLMs, SelectLLM first divides the entire dataset into several small subsets to construct the input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving the overall dataset structure. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these constructed queries into the LLM, and SelectLLM selects a few instructions that are expected to be the most helpful in fine-tuning models.
@@ 2 Related Work
One line of work focuses on density to include diverse instructions; for example, Chen et al. (2023a) map the unlabelled instructions into an embedding space and use K-means clustering (Hartigan and Wong, 1979) and the K-center greedy algorithm (Sener and Savarese, 2018) for the selection. Another line of work focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) measure uncertainty by observing how LLM-generated responses vary with changes in the input instructions. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to help train the model; from LLM’s capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. Meanwhile, a recent study by (Chen et al., 2023b) also prompts LLMs to construct a high-quality instruction dataset. However, our work differs as we focus on selecting unlabelled instructions while (Chen et al., 2023b) focuses on filtering out low-quality labeled instructions.
@@ Appendix D More Analyses with Chosen Instructions
Methods Rouge-L (F1) Cosine Similarity Dolly-15k 0.292 0.689 SelectLLM-3k 0.289 0.686
Methods Rouge-L (F1) Cosine Similarity Dolly-15k 0.292 0.689 SelectLLM-3k 0.289 0.686

### 05 Mar 2024 v4

@@ 1 Introduction
As it is relatively easy to access unlabelled instructions or texts from various sources, one promising solution to address these challenges is developing a way to select important unlabelled instructions for annotation, similar to active learning (Settles, 2009); from this, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. However, existing selection methods present limitations. Active learning-based algorithms focusing on density in the embedding space (Sener and Savarese, 2018) often overlook the complexity of instruction samples. Alternatively, output-based approaches that assess uncertainty in model predictions (Kung et al., 2023) grapple with high computational costs and the challenges posed by synthetic label quality (Figure 1). In light of these challenges, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we investigate whether LLMs can select effective unlabelled instructions, leveraging their vast knowledge base to discern the complexity and utility of each instruction without generating inferences.
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. At a high level, we use LLMs to estimate the usefulness and impact of each instruction without corresponding labels; through initial experiments presented in Figure 3, we first verify that LLMs possess such capability. To improve instruction selection via LLMs, SelectLLM first divides the dataset into several small subsets to construct input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving overall dataset structure. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these queries into the LLM, and SelectLLM selects a few instructions expected to be most helpful in fine-tuning models.
@@ 2 Related Work
One line of work focuses on density to include diverse instructions; for example, Chen et al. (2023a) map the unlabelled instructions into an embedding space and use K-means clustering (Hartigan and Wong, 1979) and the K-center greedy algorithm (Sener and Savarese, 2018) for the selection. The idea of clustering to select instruction data is also adopted in a vision-language domain (Wei et al., 2023). Another line of work focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) measure uncertainty by observing how LLM-generated responses vary with changes in the input instructions. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to help train the model; from LLM’s capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. Meanwhile, a recent study by (Chen et al., 2023b) also prompts LLMs to construct a high-quality instruction dataset. However, our work differs as we focus on selecting unlabelled instructions while (Chen et al., 2023b) focuses on filtering out low-quality labeled ones.
@@ 5 Conclusion
Acknowledgements
This project has been generously supported by funding from Cisco. We thank the Cisco team and Minnesota NLP group members for their invaluable feedback on both the project and this draft.
@@ Appendix C Verifying Capability of LLMs for Selecting Unlabelled Instruction
Methods Rouge-L (F1) Cosine Similarity Dolly-14k 0.292 0.689 SelectLLM-3k 0.289 0.686
Methods Rouge-L (F1) Cosine Similarity CleanedAlpaca-51k 0.356 0.755 SelectLLM-3k 0.301 0.707